Construct the cumulative frequency distribution for the given data.

Daily Low

​(degrees°​F)

Frequency

​35-39

3

​40-44

2

​45-49

6

​50-54

15

​55-59

5

​60-64

5

​65-69

1

Construct the cumulative frequency distribution.

Daily Low Temperature

​(degrees°​F)

Cumulative

Frequency

Less than 40

Less than 45

Less than 50

Less than 5

Less than 60

Less than 65

Less than 70
Answer #1

Solution

We are given frequency distribution as below

Daily Low

​(degrees°​F)

Frequency

​35-39

3

​40-44

2

​45-49

6

​50-54

15

​55-59

5

​60-64

5

​65-69

1

We have to find cumulative frequency distribution. Less than Cumulative frequency gives the number of observations less than or equal to upper limit of each class. To find cumulative frequency distribution we add the frequencies from top to bottom. Thus we get

Daily low temperature calculation cumulative frequency
Less than 40 3= 3
Less than 45 3+2=5 5
Less than 50 5+6=11 11
Less than 55 11+15=26 26
Less than 60 26+5=31 31
Less than 65 31+5=36 36
Less than 70 36+1=37 37